### v2

# Do More Suspicious Transaction Reports Lead to More Convictions for Money Laundering? Empirical Evidence from the European Union

Abstract
Almost all countries in the world require banks to report suspicious transactions to national authorities. The reports are known as suspicious transaction or activity reports (we use the former term) and are intended to help authorities detect and prosecute money laundering. In this paper, we investigate the relationship between suspicious transaction reports and convictions for money laundering in the European Union. We use publicly available data from Europol, the World Bank, the International Monetary Fund, and the European Sourcebook of Crime and Criminal Justice Statistics. To analyze the data, we employ a log-transformation and fit pooled (i.e., ordinary least squares) and fixed effects regression models. The fixed effects models, in particular, allow us to control for unobserved country-specific confounders (e.g., different laws regarding when and how reports should be filed). Initial results indicate that the number of suspicious transaction reports and convictions for money laundering in a country follow a sub-linear power law. Thus, while more reports may lead to more convictions, their marginal effect decreases with their amount. The relationship is robust to control variables such as the size of shadow economies and police forces. However, when we include time as a control, the relationship disappears in the fixed effects models. This suggests that the relationship is spurious rather than causal, driven by cross-country differences and a common time trend. In turn, a country cannot, ceteris paribus and with statistical confidence, expect that an increase in suspicious transaction reports will drive an increase in convictions.
Our results have important implications for international anti-money laundering efforts and policies. At best, the number of suspicious transaction reports and convictions for money laundering in a country appear to follow a sub-linear power law. At worst, there appears to be no or only a very weak (and insignificant) relationship between the number of suspicious transaction reports and convictions for money laundering in a country. Either way, our results suggest that financial regulators and supervisors should focus less on the number of suspicious transaction reports filed in a given country. Instead, they should focus more on how reports are utilized differently across countries and over time.
1 Introduction
A meta-analysis published by the United Nations Office on Drugs and Crime [61] suggests that money laundering amounts to 2.7% of the world’s combined gross domestic product (GDP). At the same time, scholars have long questioned the effectiveness of global anti-money laundering (AML) efforts [63, 58, 56, 62]. Such efforts are guided by the Financial Action Task Force (FATF), an intergovernmental organization founded in 1989. Among other things, FATF recommends that entities such as banks should be required to report suspicious transactions to national financial intelligence units [50]. The reports are known as suspicious transaction or activity reports (we use the former term) and are intended to help authorities detect and prosecute money laundering. In this paper, we investigate the relationship between suspicious transaction reports (STRs) and convictions for money laundering in the European Union (EU). Following the EU’s 4th AML directive, Member States must have “effective” AML systems and, among other things, require obligated entities to file STRs [45]. More or less implicitly, the EU directive and FATF’s AML recommendations thus assume that STRs are an effective tool to combat money laundering. This is also evident from the fact that STR counts are used as a measure of success in international AML evaluations [49, 48]. Our article explores this idea, asking if “more STRs lead to more convictions for money laundering?”
Motivated by the observation that different data sources may be inconsistent, we use STR data from a single authoritative source; [46]. The data details the number of STRs filed in EU countries between 2006 and 2014. We couple the data with statistics and estimates from the World Bank, the International Monetary Fund, and the European Sourcebook of Crime and Criminal Justice Statistics. The result is an imbalanced panel dataset with two dimensions: country and year. To model the data, we take inspiration from the Cobb-Douglas production function [42], one of the most influential production functions in economics. The function uses a power law to model the relationship between inputs (e.g., capital and labor) and output (e.g., GDP). In turn, we use a log-transformation to linearize our data and fit pooled (i.e., ordinary least squares) and fixed effects regression models. The fixed effects regression models, in particular, allow us to control for unobserved country-specific confounders, including different laws regarding when and how STRs should be filed. For illustration purposes, Figure 1 depicts our most basic pooled regression model.
Fig. 1: Illustration of our most basic pooled regression model (P1). $STR_{(c,t)}$ denotes the number of STRs filed in country $c$ in year $t$ per 100,000 capita. $CON_{(c,t+1:3)}$ denotes the average number of money laundering convictions per year in country $c$ up to three years after $t$ per 100,000 capita. We measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
Initial results indicate that the number of STRs and convictions for money laundering in a country follow a sub-linear power law, implying that while more STRs may lead to more convictions, their marginal effect decreases with their amount. For robustness, we do extensive model checks. This includes using control variables such as a country’s population size, the estimated size of its shadow economy, and the size of its police force. The relationship between STRs and convictions is robust to our controls both in the pooled and fixed effects regression models. However, when we include either a linear time trend (i.e., “time” as a regressor) or time fixed effects, the significance of the relationship disappears in the fixed effects models. This suggests that the relationship is spurious rather than causal, driven by cross-country differences and a common time trend. Thus, a country cannot, ceteris paribus and with statistical confidence, expect that an increase in STRs will drive an increase in convictions. Instead, one or more latent, country- and time-specific variables may be driving, or at least regulating, the relationship between STRs and convictions for money laundering.
Our results have important implications for international AML efforts and policies. At best, the number of STRs and convictions for money laundering in a country appears to follow a sub-linear power law. At worst, there appears to be no or only a very weak (and insignificant) relationship between the number of STRs and convictions for money laundering in a country. This would be a cause for serious concern about the effectiveness of current AML efforts and policies. Either way, our results suggest that financial regulators and supervisors should focus less on increasing the number of STRs filed in a given country. Instead, they should focus more on how STRs are utilized differently across countries and over time.
The rest of our paper is organized as follows. Section 2 reviews related work. Section 3 presents our data. Section 4 presents our methodology. Section 5 presents our results. Finally, Section 6 contains a conclusion and discussion.
2 Related Literature
Scholars have long questioned the effectiveness of global AML efforts and policies [63, 58, 56, 62]. [62] calls AML “the world’s least effective policy experiment,” arguing that less than 0.1% of criminal proceeds are seized while AML costs vastly exceed assets recovered. The author partly ascribes the failure of the global AML system to an excessive focus on activity metrics (e.g., the number of STRs filed) over outcomes (e.g., the number of convictions achieved or reduced crime rates). Somewhat more diplomatic, [58] argue that evaluation of AML efforts is hindered by an absence of reliable data. Indeed, the authors note that AML evaluations and assessments (including those by FATF) mostly rely on rhetorical arguments and narrative judgments, providing little to no evidence on effectiveness. This fits with a broader pattern, noted by [57], whereby “governments and intergovernmental organizations have spent almost nothing on public research on these issues, while regular costly (…) international meetings service the global fight against this ill-understood phenomenon and to ‘evaluate’ these efforts.” In an extensive essay and review, [56] note that “the universal message from the modest research that has been conducted is that the proportion of SARs [a version of STRs] in high-reporting jurisdictions that are actually seriously followed up is low, though the extent to which this is inherent or merely resource-constrained remains unclear.” In a broader perspective, [56] note that “how well the [AML] system works in suppressing crimes and preventing terrorist acts is entirely a matter of speculation.”
In the absence of empirical studies, the literature is heavily influenced by a theoretical model from [66]. The model shows that when financial institutions face significant fines for failing to file “enough” STRs, they may respond by filing large volumes of low-value STRs. [^2] From conversations with AML officers at multiple banks, we are aware that the officers colloquially use the terms “defensive reports” or “cover-my-ass reports” to describe such STRs; terms we believe capture the essence of the crying-wolf phenomenon. This behavior leads to an overload of low-quality information being sent to financial intelligence units (a “crying wolf” effect), reducing the value of STRs. The model’s predictions align with empirical evidence from the United States (US), where STR volumes have increased without an increase in convictions. The model is also backed up by work from [60], finding that “data suggest that Italian bankers are likely to cry wolf.”
The context of AML regimes may affect the utility of STRs. [67] compare AML regimes in the US and the Netherlands. After introducing a risk-based approach in the 1990s and 2000s, the US saw an increase in reports but a decrease in convictions for money laundering (consistent with the “crying wolf” phenomenon). Meanwhile, the Netherlands saw a decrease in reports but an increase in convictions. The authors argue that this is related to different legal traditions; the US being adversarial and the Dutch being cooperative. We note that the EU introduced its 4th AML Directive in 2015, promoting a “risk-based approach” to AML [45]. While data is scarce and fragmented, the directive does not appear to have caused a general decrease in STRs; see, for example, [40, 39, 52]. The organization of financial intelligence units may also affect the utility of STRs. [41] consider factors such as financial and informative powers, investigative authority, independence, and accountability, finding substantial differences across financial intelligence units. Finally, the type of criminal activity taking place in a country may naturally affect the utility of STRs. As noted by [61], it is, for example, likely that “the geographic concentration of cocaine trafficking activities in the Western hemisphere” means that the banking sector (typically filing more STRs than other sectors of the economy) plays a larger role in money-laundering in the Western hemisphere. In general, however, very little is known about money laundering, its predicate crimes, or the effectiveness of AML regimes [56, 58, 51].
3 Data
To build our dataset, we combine publicly available statistics and estimates from Europol, the World Bank, the International Monetary Fund, and the European Sourcebook of Crime and Criminal Justice Statistics. The result is an imbalanced panel dataset with two dimensions: country and year. Our dataset contains information on 28 EU countries between 2006 and 2016. The primary variables in our dataset are:
  • $STR_{(c,t)}$: the number of STRs filed in country $c$ in year $t$,
  • $POP_{(c,t)}$: the population size of country $c$ in year $t$.
  • $GDP_{(c,t)}$: the GDP of country $c$ in year $t$ in constant 2015 USD,
  • $SHW_{(c,t)}$: the estimated size of the shadow economy of country $c$ in year $t$ as a percentage of GDP,
  • $POL_{(c,t)}$: the number of police officers employed in country $c$ in year $t$, and
  • $CON_{(c,t)}$: the number of people convicted for money laundering in country $c$ in year $t$.
By multiplying $SHW_{(c,t)}$ and $GDP_{(c,t)}$, we construct another variable:
  • $SGDP_{(c,t)}$: the estimated size of the shadow economy of country $c$ in year $t$ in constant 2015 USD.
The idea behind $SGDP_{(c,t)}$ (shorthand for “Shadow GDP”) is to express the size of a country’s shadow economy in monetary terms rather than as a percentage of GDP (something more natural for the models we consider in Section 4). In our models, tables, and graphs, we measure $GDP_{(c,t)}$ and $SGDP_{(c,t)}$ per capita (simply dividing by $POP_{(c,t)}$). Meanwhile, we measure $STR_{(c,t)},POL_{(c,t)},$ and $CON_{(c,t)}$ per 100,000 capita (dividing by $POP_{(c,t)}$ and multiplying by 100,000). To ease notation later, we measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
$STR_{(c,t)}$ $GDP_{(c,t)}$ $SHW_{(c,t)}$ $SGDP_{(c,t)}$ $POL_{(c,t)}$ $CON_{(c,t)}$ $POP_{(c,t)}$ $t$
Mean 231.47 29,774.35 22.98 5,624.12 343.01 0.99 18,022,396.17 5.00
Std. dev. 430.23 20,825.57 8.07 2,294.32 104.85 1.31 22,840,247.64 3.17
Minimum 3.56 5,629.80 9.10 2,037.99 58.18 0.00 405,308.00 0.00
Median 41.69 21,507.64 21.40 5,382.46 340.16 0.34 8,908,586.50 5.00
Maximum 2,364.41 112,417.88 38.50 10,618.44 570.67 6.45 82,376,451.00 10.00
Table 1: Summary statistics for our dataset, including countries and years with missing observations. “Std. dev.” denotes standard deviation. We use $t$ to denote the year of an observation, measured relative to 2006. Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on. Note that $GDP_{(c,t)}$ and $SGDP_{(c,t)}$ are measured per capita. Meanwhile, $STR_{(c,t)},POL_{(c,t)},$ and $CON_{(c,t)}$ are measured per 100,000 capita.
Table 1 presents summary statistics for our dataset. Note that there is large variation in $STR_{(c,t)}$, having a standard deviation of 430.23 (substantially more than its mean of 231.47). Table 2 presents correlation coefficients for our dataset (calculated by comparing observations from the same year, i.e., without a lag). Note that $GDP_{(c,t)},SHW_{(c,t)}$ and $SGDP_{(c,t)}$ (quite naturally) are strongly correlated. While $STR_{(c,t)}$ and $t$ only appear to be weakly correlated, an inspection of our dataset reveals a highly non-linear increase in STRs within countries over time. Indeed, the number of STRs per 100,000 capita increased by almost 240% on average, taken across all countries, between 2006 and 2014 (with a median increase of approximately 140%).
The following subsections present each of our data sources and variables in detail. Subsection 3.1 presents our $STR_{(c,t)}$ statistics from Europol. Subsection 3.2 presents our $GDP_{(c,t)}$ and $POP_{(c,t)}$ statistics from the World Bank. Subsection 3.3 presents our $SHW_{(c,t)}$ estimates from the International Monetary Fund. Finally, Subsection 3.4 presents our $POL_{(c,t)}$ and $CON_{(c,t)}$ statistics from the European Sourcebook of Crime and Criminal Justice Statistics. Our dataset is available at https://github.com/TuffvesonJensen/STR2CON.
$STR_{(c,t)}$ $GDP_{(c,t)}$ $SHW_{(c,t)}$ $SGDP_{(c,t)}$ $POL_{(c,t)}$ $CON_{(c,t)}$ $POP_{(c,t)}$ $t$
$STR_{(c,t)}$ -1.00 -0.26 -0.17 -0.12 -0.24 -0.45 -0.16 -0.07
$GDP_{(c,t)}$ -0.26 -1.00 -0.76 -0.80 -0.50 -0.23 -0.07 -0.02
$SHW_{(c,t)}$ -0.17 -0.76 -1.00 -0.43 -0.38 -0.10 -0.28 -0.02
$SGDP_{(c,t)}$ -0.12 -0.80 -0.43 -1.00 -0.30 -0.35 -0.03 -0.01
$POL_{(c,t)}$ -0.24 -0.50 -0.38 -0.30 -1.00 -0.06 -0.02 -0.11
$CON_{(c,t)}$ -0.45 -0.23 -0.10 -0.35 -0.06 -1.00 -0.33 -0.34
$POP_{(c,t)}$ -0.16 -0.07 -0.28 -0.03 -0.02 -0.33 -1.00 -0.01
t -0.07 -0.02 -0.02 -0.01 -0.11 -0.34 -0.01 -1.00
Table 2: Correlation coefficients (without any time lag) over all countries and years in our dataset, including those with missing observations. We use $t$ to denote the year of an observation, measured relative to 2006. Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
3.1 Statistics on Suspicious Transaction Reports From Europol
Different countries have different laws regarding when and how reports about suspicious transactions should be filed. In turn, the definition of (and name given to) an STR is open to debate, making it difficult to obtain comparable statistics from national authorities. Reports are sometimes denoted as suspicious activity reports (SARs), unusual transaction reports (UTRs), or notifications. The Netherlands, for example, operates a UTR regime [46], Denmark primarily operates an STR regime [47], and Poland operates a regime with multiple types of SARs and notifications [59]. Motivated by the observation that different sources may be inconsistent, we use STR data from a single authoritative source; [46]. [^3] We note that even statistics published in the annual reports of a single financial intelligence unit can be inconsistent from year to year (we omit examples in order not to single out individual authors or financial intelligence units). In turn, we follow [46] and define an STR as any report received by a national financial intelligence unit from an obligated entity, whether called an STR, SAR, UTR, or notification. [46] lists the number of STRs filed in each EU country from 2006 to 2014, yielding our variable:
  • $STR_{(c,t)}$: the number of STRs filed in country $c$ in year $t$.
Given 28 countries and a data collection period of 9 years, we have 252 observations of $STR_{(c,t)}$. In our models, tables, and graphs, we measure $STR_{(c,t)}$ per 100,000 capita.
3.2 Statistics on GDP and Population Sizes From the World Bank
We use the online database of the [70] to obtain the following variables:
  • $GDP_{(c,t)}$: the GDP (World Bank indicator code NY.GDP.MKTP.KD) of country $c$ in year $t$ in constant 2015 USD and
  • $POP_{(c,t)}$: the population size (World Bank indicator code SP.POP.TOTL) of country $c$ in year $t$.
We collect both variables from 2006 (the start of our STR data) to 2016 (the end of our conviction data). Given 28 countries and a data collection period of 11 years, we have 308 observations of each variable. In our models, tables, and graphs, we measure $GDP_{(c,t)}$ per capita.
3.3 Estimates of the Shadow Economy from the International Monetary Fund
We use shadow economy estimates published by the International Monetary Fund [55]. The estimates build on work by [54] and are derived from a multiple indicators, multiple causes model. The model infers the size of a country’s shadow economy from macroeconomic indicators, including productivity, tax revenue, trade volume, labor force participation, and investment; see [55] and [54]. Thus, we record the following variable from [55]:
  • $SHW_{(c,t)}$: the estimated size of the shadow economy of country $c$ in year $t$ as a percentage of GDP.
We collect the variable from 2006 (the start of our STR data) to 2016 (the end of our conviction data). Given 28 countries and a data collection period of 11 years, we have 308 potential observations of $SHW_{(c,t)}$. However, [55] does not report estimates for Malta. In turn, we only have 297 observations.
We stress that “the shadow economy” and “money laundering” are distinct but related concepts. In particular, the shadow economy is defined as economic activity hidden from authorities [55, 54, 53, 69]. This will largely include activities that, in and of themselves, are legal. Furthermore, such activities do not necessarily involve money laundering; just as all criminal activities do not require money laundering. [^4] Imagine, for example, a day laborer who is paid in cash. The laborer’s work may, in and of itself, be legal. If the laborer fails to declare their income, it constitutes tax evasion. However, if the laborer simply uses their cash salary (including what should have been paid in taxes) to pay for daily expenses such as groceries, they may not be involved in any kind of money laundering.
3.4 Statistics on Police Forces and Convictions for Money Laundering From the European Sourcebook of Crime and Criminal Justice Statistics
We use the European Sourcebook of Crime and Criminal Justice Statistics [37, 36, 38] as our source of police and conviction statistics. [^5] We extract and use raw statistics from the Sourcebook’s databases, available on the Sourcebook’s website [37]. These statistics may at times differ from those published in the print versions of the Sourcebook. The Sourcebook collects crime and criminal justice statistics from national correspondents. The 6th edition of the Sourcebook, for example, contains statistics from the Danish Ministry of Justice, the Latvian Statistical Bureau, and numerous academics from across Europe. Notably, the Sourcebook aims to collect standardized statistics; a prerequisite for comparing statistics across countries. [^6] As an alternative data source, we considered using mutual evaluation reports from FATF. However, statistics in these are not standardized. For example, [47] combines statistics on money laundering and stolen goods. While the Sourcebook aims to collect standardized statistics, these may deviate. In this work, we simply use the statistics as given in the Sourcebook. The Sourcebook explicitly defines money laundering as “specific financial transactions to conceal the identity, source, and/or destination of money or non-monetary property deriving from criminal activities.” We collect the following variables from the Sourcebook:
  • $POL_{(c,t)}$: the number of police officers employed (Sourcebook indicator code T14OP) in country $c$ in year $t$ and
  • $CON_{(c,t)}$: the number of people convicted for money laundering (Sourcebook indicator code T31ML) in country $c$ in year $t$.
We collect both variables from 2006 (the start of our STR data) to 2016 (two years after the end of our STR data). The latter is motivated by the fact that the 6th edition of the Sourcebook [37], the last to be fully published, covers the years 2011-2016.777While a preliminary 7th edition is available, we do not consider it as its underlying data is unpublished at the time of writing. Multiple previous editions of the Sourcebook have been published. A 5th edition [36] covers the years 2007-2011 while a 4th edition [38] covers the years 2003-2007. Note that each edition of the Sourcebook overlaps slightly with its predecessor. This can cause contradictions. For example, the 5th edition reports zero convictions in Slovenia in 2011. Meanwhile, the 6th edition reports one conviction in Slovenia in 2011. Whenever contradictions arise, we use data from the most recent edition. [^8] For a given country and year, one edition might report a missing observation. Meanwhile, another edition might report an observation. In this case, we use the available observation, regardless of whether it comes from the most recent edition. We note that there are substantial “jumps” in $POL_{(c,t)}$ for some countries between different editions of the Sourcebook. This might indicate that the statistics in the Sourcebook are not fully standardized. Each edition of the Sourcebook also contains missing observations. Thus, our panel data becomes imbalanced. Given 28 countries and a data collection period of 11 years, we have 308 possible observations. However, we only have 245 observations of $POL_{(c,t)}$ and 179 observations of $CON_{(c,t)}$. [^9] Data on the United Kingdom (UK), a member of the EU until 2020, is somewhat ambiguous in the Sourcebook as statistics are provided on “UK: England & Wales,” “UK: Northern Ireland,” and “UK: Scotland.” We simply sum these. In Section 5, we run extensive robustness checks, sequentially excluding countries from our models to check if our results still hold. In our models, tables, and graphs, we measure $POL_{(c,t)}$ and $CON_{(c,t)}$ per 100,000 capita.
Figures 2 and 3 illustrate the number of observations of $CON_{(c,t)}$ per country and year in our dataset. As can be seen, there are some countries for which we have no observations. Still, our dataset covers Northern, Southern, Eastern, and Western Europe. As can also be seen, the number of observations increases towards the end of our data collection period. We believe this is due to improvements in the way data was collected in the Sourcebook; not interaction effects between STRs and convictions. Interestingly, the Sourcebook also contains statistics on convictions for some predicate crimes to money laundering; we consider these in Appendix C.
Fig. 2: Observations of $CON_{(c,t)}$ per country $c$. We have 179 observations in total. There are five countries for which we have no observations: France, Greece, Latvia, Luxembourg, and Malta.
Fig. 3: Observations of $CON_{(c,t)}$ per year $t$. Note that $t$ is measured relative to 2006 (the first year of our data). We have 179 observations in total. The number of observations increases towards the end of our data collection period.
4 Methodology
We use pooled and fixed effects regression models to investigate if an increase in STRs leads to an increase in convictions for money laundering. Both types of models are standard in econometric analysis (see, for example, [68, 43]). Pooled regression treats panel data as cross-sectional (i.e., as though all observations are independent; corresponding to ordinary least squares regression). While unrealistic, this provides a useful starting point for our analysis. Fixed effects regression, by contrast, uses entity-specific terms to capture and account for unobserved entity-specific confounders.
Cases about money laundering can take multiple years to investigate and go through criminal justice systems. Motivated by this, we use a time lag to model the relationship between STRs and convictions for money laundering. To be specific, we consider the impact STRs have on convictions up to and including five years after they have been filed. [^10] According to the [44], court cases on money laundering in most EU countries on average took around 300 days or less in 2014. In some countries, however, the average was over 600 days. Allowing additional time for investigation, we believe five years is a reasonable time-frame to consider. We stress that we do not claim STRs are useful only within five years of being filed. Rather, we simply claim that the principal effect of an increase in STRs should be visible within five years. Recall that $STR_{(c,t)}$ denotes the number of STRs in country $c$ in year $t$ and that $CON_{(c,t)}$ denotes the number of convictions in country $c$ in year $t$. Now, let $CON_{(c,t+j)}$ denote the number of convictions in country $c$ in year $t+j$, i.e., $j$ years after $t$, and let
$CON_{(c,t+1:H)}=\frac{1}{H}\sum_{j=1}^{H}CON_{(c,t+j)}$ (1)
denote the average number of convictions in country $c$ per year up to and including $H$ years after $t$. We measure $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ per 100,000 capita using $POP_{(c,t)}$. [^11] This is based on the idea that $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ are determined by $STR_{(c,t)}$, i.e., the number of STRs in year $t$, not years t+1,,t+Ht+1,...,t+H. For robustness, we also tried measuring $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ using population counts in year $t+j$ and the average population count over years t+1,,t+Ht+1,...,t+H. This does not materially change our conclusions. We now consider and model:
  1. 1.
    the relationship between $STR_{(c,t)}$ and $CON_{(c,t+j)}$ for $j=1,2,3,4,5$ and
  2. 2.
    the relationship between $STR_{(c,t)}$ and $CON_{(c,t+1:H)}$ for $H=3$ and H=5H=5.121212Note that using $CON_{(c,t+1:H)}$ as a dependent variable naturally introduces serial correlation between observations.
In the following subsections, we present our pooled and fixed effects regression models. Throughout the subsections, we use $CON_{(c,t+j)}$ as the dependent variable. However, our models for $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ are identical; the reader may simply substitute the dependent variable. Subsection 4.1 presents our pooled regression model. Subsection 4.2 presents our fixed effects regression model. Note that the number of observations used to fit each model naturally changes depending on the time period ($j$ and $H$) being considered. Throughout Section 5, we always specify the number of observations used to fit each model. Our implementation relies on the statsmodels [64] and linearmodels [65] Python libraries. A replication package is available at https://github.com/TuffvesonJensen/STR2CON.
4.1 Pooled Regression
Inspired by the Cobb-Douglas production function, we may model $STR_{(c,t)}$ and $CON_{(c,t+j)}$ as
$CON_{(c,t+j)}=A\times STR_{(c,t)}^{\beta}$ (2)
where $A\in\mathbb{R}_{>0}$ measures a “scale of production” and $\beta\in\mathbb{R}$ measures how convictions respond to a change in STRs. [^13] As an alternative to the power-law formulation, we also considered a simple linear formulation; see Appendix B. To fit the model, we use a log-transformation, yielding an ordinary least squares regression
$\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})$ (3)
where $\alpha=\log_{10}(A)$ is an intercept and $\beta$ is a regression coefficient. [^14] Our dataset has 9 observations of $CON_{(c,t)}=0$ which cannot undergo a log-transformation. When we fit our models, we dynamically drop such observations. In our data, no observations of $CON_{(c,t+1:H)}$ are equal to zero (being averages over multiple years). Thus, no observations of $CON_{(c,t+1:H)}$ are dropped. As an alternative to “drop-handling,” we tried adding $0.01$ to all $CON_{(c,t+j)}$ observations (after taking per capita rates but before log-transforming). This does not change our material conclusions. The model pools all our observations, treating them as cross-sectional. In particular, this means that the model ignores unobserved country-specific confounders (e.g., different laws regarding when and how STRs should be filed). This may bias our results.
Motivated by the observation that we have per capita ratios on both sides of equation (3) (note that both $CON_{(c,t+j)}$ and $STR_{(c,t)}$ are measured per 100,000 capita) and that a country’s population size might have a stand-alone effect on conviction rates, we include $log_{10}(POP_{(c,t)})$ as a control variable in the model in (3). [^15] Algebraically, $\log_{10}(y/n)=\alpha+\beta\,\log_{10}(x/n)+u\iff\log_{10}y=\alpha+\beta\,\log_{10}(x)+(1-\beta)\,\log_{10}(n)+u$. Thus, omitting $\log_{10}(n)$ (or, in our case, $\log_{10}(POP_{(c,t)})$) as a control variable imposes a (hidden) restriction that its coefficient must equal $1-\beta$. Note, also, that measuring variables per (100,000) capita shifts the intercept of our models. Furthermore, we sequentially log-transform and add additional control variables for robustness. In particular, the amount of money laundering in a country could, in and of itself, drive both STRs and convictions. Imagine, for example, that banks consistently file STRs on 10% of all money launderers and that 5% of all launderers are convicted regardless of the number of STRs filed. In this case, an increase in launderers “mechanically” leads to more STRs and convictions, creating a classic case of omitted variable bias (i.e., the illusion of a causal relationship). To address this, we would like to control for the amount of money laundering in a country; something we, for obvious reasons, do not have data on. Instead, we add $\log_{10}(SGDP_{(c,t)})$ as a control variable. This is motivated by the idea that $SGDP_{(c,t)}$, estimating the size of a country’s shadow economy, is a reasonable proxy for money laundering activities when we consider comparable countries. Indeed, we find it is reasonable to suspect that a larger shadow economy, all else being equal, might increase demand for money laundering. Furthermore, a larger shadow economy might also, in itself, influence the number of STRs being filed. The covert nature of money laundering means that we cannot directly investigate the relationship between a country’s shadow economy and the amount of money laundering in said country (if we knew the amount of money laundering in a country, we would not need a proxy for it). Note, however, that we only use $\log_{10}(SGDP_{(c,t)})$ as a control variable. [^16] The reader may be concerned that $SGDP_{(c,t)}$ builds on an estimate from a model and is not directly observed. We agree with this concern but note that illicit financial flows (by their nature) cannot be observed. We also note that GDP figures, commonly used in econometric analysis, come from statistical models (though they might be very well-developed). We also add $\log_{10}(POL_{(c,t)})$ as a control variable. This is motivated by the idea that convictions for money laundering may depend on the number of police officers in a country. Indeed, in the classic Cobb-Douglas setup, one might think of $STR_{(c,t)}$ as capital and $POL_{(c,t)}$ as labor. Finally, we also try to control for time using, respectively, (i) $t$ (i.e., time) as a regressor (assuming a log-linear time trend) and (ii) time dummy variables. [^17] Our model in (i) may, assuming no other control variables, be specified as $\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})+\delta\times t$ where $\delta\in\mathbb{R}$ is a regression coefficient. Similarly, our model in (ii) may be specified as $\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})+\sum_{k=1}^{K}\delta_{(k)}\mathbf{1}_{\{t=k\}}$ where $K$ denotes the last year $t$ for which we have an observation of $STR_{(c,t)}$ to model and $\delta_{(1)},\dots,\delta_{(K)}$ are regression coefficients. For numerical stability, $t$ is measured relative to 2006. We dynamically drop observations if one of our control variables is missing. Standard errors are clustered at the unit (i.e., country) level as implemented by statsmodels [64].
4.2 Fixed Effects Regression
Inspired by the Cobb-Douglas production function, we may model $STR_{(c,t)}$ and $CON_{(c,t+j)}$ as
$CON_{(c,t+j)}=A_{(c)}\times STR_{(c,t)}^{\beta}$ (4)
where $A_{(c)}\in\mathbb{R}_{>0}$ is a country-specific fixed effect and $\beta\in\mathbb{R}$ measures how convictions (in general) respond to a change in STRs. Notably, $A_{(c)}$ measures the “scale of production” in country $c$, i.e., how “good” a particular country $c$ is at converting STRs into convictions. As discussed in Section 2, there are many reasons why countries may have different $A_{(c)}$ values (including different laws regarding when and how STRs should be filed). To fit the model, we use a log-transformation, yielding a fixed effects regression
$\log_{10}(CON_{(c,t+j)})=\alpha_{(c)}+\beta\times\log_{10}(STR_{(c,t)})$ (5)
where $\alpha_{(c)}=\log_{10}(A_{(c)})$ is a country-specific fixed effect and $\beta$ is a regression coefficient. For estimation, one could use country-specific dummy variables to represent $\alpha_{(c)}$. However, for numerical reasons, it is standard to demean observations at the unit level; see, e.g., [43]. The model naturally controls for unobserved country-specific confounders that are constant over time.
As in Subsection 4.1, we add $\log_{10}(POP_{(c,t)})$, $\log_{10}(SGDP_{(c,t)})$, and $\log_{10}(POL_{(c,t)})$ as control variables to the model in (5). Furthermore, we try to control for time. To this end, we, respectively, include (i) $t$ (i.e., time) as a regressor (assuming a log-linear time trend) and (ii) time fixed effects. [^18] Our model in (i) may, assuming no other control variables, be specified as $\log_{10}(CON_{(c,t+j)})=\alpha_{(c)}+\beta\times\log_{10}(STR_{(c,t)})+\delta\times t$ where $\delta\in\mathbb{R}$ is a regression coefficient. Similarly, our model in (ii) may be specified as $\log_{10}(CON_{(c,t+j)})=\alpha_{(c)}+\beta\times\log_{10}(STR_{(c,t)})+\delta_{(t)}$ where $\delta_{(t)}\in\mathbb{R}$ is a time fixed effect. For numerical stability, $t$ is measured relative to 2006. For the model in (ii), standard errors are clustered both at the unit and time level as implemented by linearmodels. Unless otherwise stated, we cluster standard errors at the unit (i.e., country) level as implemented by linearmodels [65].
5 Results
We first consider results from our pooled regression models, treating all observations as independent. Next, we consider results from our fixed effects regression models, allowing more robust inference by controlling for unobserved country-specific confounders. Subsection 5.1 presents our pooled regression results. Subsection 5.2 presents our fixed effects regression results. Throughout both subsections, we test the robustness of our results by varying dependent variables, sequentially excluding countries, and including control variables and time effects. We generally round results to two decimals and evaluate statistical significance at a $0.05$ $p$-level. The European Sourcebook of Crime and Criminal Justice Statistics also contains statistics on convictions for some predicate crimes to money laundering: we present results using these in Appendix C.
5.1 Pooled Regression Results
Table 3 presents our pooled regression results, considering multiple model specifications, named P1 to P4, and multiple dependent variables as explained in Section 4. In model P1, we only use $\log_{10}(STR_{(c,t)})$ as a regressor. In model P2, we include $\log_{10}(POP_{(c,t)})$ as a control variable. In model P3, we include $\log_{10}(SGDP_{(c,t)})$ as a control variable. Finally, in model P4, we include $\log_{10}(POL_{(c,t)})$ as a control variable. For illustration purposes, Figure 1 depicts model P1. A Q-Q plot of residuals is displayed in Appendix A.
Model ID $n$ $R^{2}$ $\alpha$ $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$
Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$
P1 141 0.28 -1.20*** (0.27) 0.47*** (0.12)
P2 141 0.39 -4.23*** (1.26) 0.49*** (0.12) 0.43* (0.20)
P3 141 0.48 -8.07*** (1.94) 0.38** (0.13) 0.42* (0.18) 1.11** (0.39)
P4 132 0.49 -12.18*** (2.44) 0.48*** (0.14) 0.41* (0.17) 1.38*** (0.35) 1.19* (0.56)
Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$
P1 147 0.29 -1.13*** (0.29) 0.47*** (0.13)
P2 147 0.42 -4.39*** (1.26) 0.49*** (0.12) 0.46* (0.20)
P3 147 0.52 -8.43*** (1.88) 0.38** (0.13) 0.44* (0.18) 1.19** (0.39)
P4 137 0.54 -12.65*** (2.20) 0.48*** (0.13) 0.43** (0.16) 1.45*** (0.32) 1.24* (0.53)
Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$
P1 135 0.27 -1.03*** (0.29) 0.45*** (0.13)
P2 135 0.39 -4.11** (1.30) 0.47*** (0.12) 0.43* (0.20)
P3 135 0.52 -8.41*** (1.92) 0.36** (0.13) 0.40* (0.18) 1.27** (0.42)
P4 128 0.58 -13.18*** (2.03) 0.49*** (0.13) 0.38* (0.15) 1.54*** (0.31) 1.48** (0.49)
Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$
P1 125 0.23 -0.90** (0.28) 0.40** (0.13)
P2 125 0.34 -3.88** (1.35) 0.42*** (0.12) 0.42* (0.21)
P3 125 0.49 -8.49*** (1.99) 0.31* (0.14) 0.38* (0.19) 1.36** (0.45)
P4 121 0.54 -12.09*** (1.93) 0.43*** (0.13) 0.35* (0.16) 1.48*** (0.34) 1.27* (0.52)
Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$
P1 113 0.20 -0.78** (0.29) 0.36** (0.13)
P2 113 0.32 -3.70** (1.39) 0.38** (0.12) 0.41 (0.21)
P3 113 0.49 -8.51*** (2.04) 0.27* (0.13) 0.37 (0.19) 1.42** (0.49)
P4 111 0.56 -12.08*** (1.99) 0.41** (0.13) 0.32* (0.16) 1.52*** (0.36) 1.34* (0.59)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
P1 122 0.29 -1.19*** (0.27) 0.48*** (0.12)
P2 122 0.44 -4.77*** (1.32) 0.48*** (0.12) 0.51* (0.20)
P3 122 0.50 -8.26*** (2.00) 0.39** (0.13) 0.50** (0.19) 0.99** (0.37)
P4 116 0.50 -10.58*** (2.77) 0.45** (0.16) 0.51** (0.18) 1.15** (0.38) 0.65 (0.65)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
P1 80 0.40 -1.07*** (0.20) 0.45*** (0.08)
P2 80 0.51 -3.71** (1.15) 0.45*** (0.09) 0.38* (0.17)
P3 80 0.62 -7.94*** (1.77) 0.36*** (0.10) 0.38* (0.15) 1.17*** (0.31)
P4 80 0.67 -11.20*** (2.05) 0.46*** (0.12) 0.39** (0.13) 1.33*** (0.24) 0.99* (0.48)
Table 3: Pooled regression results. The table contains results for multiple model specifications, named P1 to P4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
All $\log_{10}(STR_{(c,t)})$ coefficients in Table 3 are statistically significant, positive, and smaller than 1; the largest being equal to $0.49$ with a standard error of $0.12$ and a 95% confidence interval of $[0.26:0.72]$. This indicates that STRs and convictions for money laundering follow a sub-linear power law. [^19] The reader may wonder why $\log_{10}(POL_{(c,t)})$ is insignificant using $\log_{10}(CON_{(c,t+1:3)})$ as the dependent variable when it is significant using $\log_{10}(CON_{(c,t+1)})$, $\log_{10}(CON_{(c,t+2)})$, or $\log_{10}(CON_{(c,t+3)})$ as the dependent variable. Note, however, that it takes three subsequent years of observations to compute $\log_{10}(CON_{(c,t+1:3)})$, why it is fitted with less and different observations.
For robustness, we sequentially exclude any one country from our dataset and re-run all models. Excluding either Austria, Belgium, Denmark, Estonia, Germany, Italy, the Netherlands, Portugal, or Slovenia causes at least one $\log_{10}(STR_{(c,t)})$ coefficient to become insignificant across models P3 and P4. As a further robustness check, we also consider all models with, respectively, (i) $t$ as a regressor (assuming a log-linear time trend) and (ii) year-specific dummy variables. Results are displayed in Tables 4 and 5. Including $t$ as a regressor generally causes a small drop in our $\log_{10}(STR_{(c,t)})$ coefficients though they mostly remain significant. When we include year-specific dummy variables, the $\log_{10}(STR_{(c,t)})$ coefficients, again, generally become a bit smaller but mostly remain significant.
Model ID $n$ $R^{2}$ $\alpha$ $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$ $t$
Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$
PT1 141 0.37 -1.46*** (0.23) 0.43*** (0.12) 0.08** (0.02)
PT2 141 0.47 -4.23*** (1.21) 0.45*** (0.11) 0.40* (0.19) 0.07** (0.02)
PT3 141 0.55 -7.96*** (1.96) 0.34** (0.13) 0.38* (0.17) 1.08** (0.41) 0.07** (0.02)
PT4 132 0.56 -11.15*** (2.35) 0.43** (0.14) 0.38* (0.16) 1.26*** (0.37) 0.97 (0.59) 0.06* (0.03)
Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$
PT1 147 0.36 -1.37*** (0.25) 0.44*** (0.12) 0.07** (0.02)
PT2 147 0.48 -4.48*** (1.21) 0.46*** (0.12) 0.44* (0.19) 0.06*** (0.02)
PT3 147 0.58 -8.52*** (1.91) 0.35** (0.13) 0.42* (0.17) 1.19** (0.41) 0.06*** (0.02)
PT4 137 0.60 -12.02*** (2.14) 0.43** (0.14) 0.42** (0.15) 1.37*** (0.35) 1.07 (0.58) 0.06** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$
PT1 135 0.32 -1.23*** (0.26) 0.42*** (0.13) 0.06** (0.02)
PT2 135 0.44 -4.22** (1.28) 0.44*** (0.12) 0.42* (0.20) 0.06** (0.02)
PT3 135 0.56 -8.51*** (1.95) 0.33* (0.14) 0.39* (0.18) 1.27** (0.44) 0.06** (0.02)
PT4 128 0.61 -12.76*** (1.97) 0.46*** (0.13) 0.38* (0.15) 1.48*** (0.33) 1.35* (0.52) 0.05* (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$
PT1 125 0.26 -1.06*** (0.27) 0.39** (0.13) 0.06** (0.02)
PT2 125 0.37 -3.95** (1.36) 0.41*** (0.12) 0.41 (0.21) 0.05** (0.02)
PT3 125 0.52 -8.57*** (2.03) 0.30* (0.14) 0.37 (0.19) 1.36** (0.46) 0.05** (0.02)
PT4 121 0.56 -11.84*** (1.91) 0.41** (0.14) 0.35* (0.16) 1.45*** (0.35) 1.18* (0.55) 0.05* (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$
PT1 113 0.21 -0.87** (0.29) 0.36** (0.13) 0.04* (0.02)
PT2 113 0.33 -3.77** (1.41) 0.37** (0.12) 0.41 (0.21) 0.04* (0.02)
PT3 113 0.51 -8.62*** (2.06) 0.26 (0.13) 0.37 (0.20) 1.43** (0.49) 0.04* (0.02)
PT4 111 0.57 -12.04*** (1.98) 0.39** (0.13) 0.32* (0.16) 1.52*** (0.37) 1.28* (0.60) 0.04* (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
PT1 122 0.39 -1.48*** (0.24) 0.44*** (0.12) 0.09** (0.03)
PT2 122 0.52 -4.84*** (1.26) 0.45*** (0.12) 0.48* (0.19) 0.08** (0.03)
PT3 122 0.58 -8.33*** (2.00) 0.36** (0.13) 0.48** (0.17) 0.99** (0.38) 0.08** (0.03)
PT4 116 0.57 -9.53*** (2.68) 0.39* (0.16) 0.48** (0.17) 1.05** (0.40) 0.37 (0.68) 0.08** (0.03)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
PT1 80 0.47 -1.28*** (0.19) 0.43*** (0.08) 0.08* (0.03)
PT2 80 0.57 -3.84*** (1.10) 0.43*** (0.08) 0.37* (0.16) 0.08** (0.03)
PT3 80 0.68 -7.91*** (1.71) 0.34*** (0.10) 0.37** (0.14) 1.13*** (0.30) 0.08** (0.03)
PT4 80 0.70 -10.33*** (2.02) 0.42*** (0.12) 0.38** (0.13) 1.26*** (0.25) 0.74 (0.50) 0.06* (0.03)
Table 4: Pooled regression results with $t$ (i.e., “time”) as a regressor. The table contains results for multiple model specifications, named PT1 to PT4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
Model ID $n$ $R^{2}$ $\alpha$ $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$ $t=1$ $t=2$ $t=3$ $t=4$ $t=5$ $t=6$ $t=7$ $t=8$ Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$ PTD1 141 0.38 -1.40*** (0.25) 0.43*** (0.12) -0.03 (0.13) 0.03 (0.16) 0.07 (0.15) 0.34* (0.17) 0.34 (0.18) 0.41* (0.19) 0.46* (0.19) 0.54** (0.17) PTD2 141 0.48 -4.16*** (1.24) 0.44*** (0.11) 0.40* (0.19) -0.06 (0.15) -0.03 (0.16) 0.06 (0.16) 0.28 (0.16) 0.28 (0.17) 0.34 (0.18) 0.39* (0.18) 0.48** (0.17) PTD3 141 0.55 -7.88*** (2.00) 0.34* (0.14) 0.39* (0.17) 1.08* (0.42) -0.09 (0.14) -0.04 (0.16) 0.05 (0.16) 0.25 (0.15) 0.24 (0.16) 0.31 (0.17) 0.37* (0.16) 0.46** (0.16) PTD4 132 0.56 -11.06*** (2.41) 0.42** (0.15) 0.38* (0.16) 1.26*** (0.38) 0.97 (0.61) -0.11 (0.14) -0.04 (0.16) 0.02 (0.16) 0.20 (0.16) 0.19 (0.17) 0.31 (0.19) 0.36 (0.19) 0.41* (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$ PTD1 147 0.38 -1.45*** (0.26) 0.44*** (0.13) 0.09 (0.11) 0.09 (0.12) 0.40* (0.17) 0.39* (0.18) 0.46* (0.18) 0.52** (0.19) 0.58** (0.18) 0.49** (0.18) PTD2 147 0.49 -4.50*** (1.20) 0.46*** (0.12) 0.44* (0.19) 0.04 (0.10) 0.11 (0.12) 0.36* (0.15) 0.35* (0.16) 0.41** (0.16) 0.47** (0.17) 0.54*** (0.16) 0.45** (0.16) PTD3 147 0.59 -8.48*** (1.93) 0.35* (0.14) 0.41* (0.17) 1.17** (0.42) 0.06 (0.11) 0.11 (0.13) 0.33* (0.15) 0.32* (0.16) 0.39* (0.16) 0.46** (0.17) 0.54** (0.16) 0.45** (0.16) PTD4 137 0.60 -11.92*** (2.21) 0.43** (0.14) 0.42** (0.16) 1.36*** (0.36) 1.05 (0.60) 0.07 (0.12) 0.08 (0.13) 0.28 (0.16) 0.26 (0.17) 0.34 (0.18) 0.46* (0.20) 0.48* (0.20) 0.44* (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$ PTD1 135 0.34 -1.33*** (0.24) 0.43*** (0.13) 0.05 (0.10) 0.31* (0.15) 0.32* (0.15) 0.40** (0.15) 0.46** (0.16) 0.53*** (0.15) 0.42** (0.16) PTD2 135 0.45 -4.32*** (1.28) 0.44*** (0.12) 0.42* (0.20) 0.10 (0.11) 0.31* (0.13) 0.33* (0.13) 0.39** (0.13) 0.46** (0.14) 0.53*** (0.13) 0.41** (0.13) PTD3 135 0.57 -8.48*** (1.97) 0.34* (0.14) 0.39* (0.18) 1.24** (0.45) 0.06 (0.11) 0.25* (0.12) 0.25* (0.12) 0.32** (0.11) 0.39** (0.13) 0.48*** (0.12) 0.39** (0.13) PTD4 128 0.62 -12.69*** (2.08) 0.46*** (0.14) 0.38* (0.16) 1.48*** (0.34) 1.34* (0.55) 0.02 (0.10) 0.15 (0.12) 0.15 (0.12) 0.23 (0.12) 0.30* (0.13) 0.39** (0.15) 0.32* (0.15) Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$ PTD1 125 0.28 -1.22*** (0.25) 0.39** (0.13) 0.28 (0.16) 0.25 (0.18) 0.37* (0.18) 0.43* (0.19) 0.50** (0.19) 0.40* (0.19) PTD2 125 0.38 -4.02** (1.40) 0.41*** (0.12) 0.40 (0.21) 0.23 (0.14) 0.20 (0.16) 0.31 (0.16) 0.37* (0.17) 0.45** (0.17) 0.34* (0.17) PTD3 125 0.53 -8.57*** (2.07) 0.30* (0.14) 0.37 (0.20) 1.35** (0.47) 0.19 (0.13) 0.14 (0.15) 0.24 (0.15) 0.32* (0.16) 0.41* (0.17) 0.33* (0.16) PTD4 121 0.57 -11.87*** (1.98) 0.41** (0.14) 0.34* (0.17) 1.46*** (0.36) 1.18* (0.57) 0.15 (0.14) 0.08 (0.18) 0.19 (0.18) 0.26 (0.18) 0.34 (0.20) 0.30 (0.20) Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$ PTD1 113 0.22 -0.86** (0.29) 0.36** (0.13) -0.02 (0.10) 0.07 (0.09) 0.16 (0.10) 0.22* (0.10) 0.13 (0.09) PTD2 113 0.34 -3.76** (1.43) 0.37** (0.12) 0.41 (0.22) -0.02 (0.09) 0.06 (0.09) 0.15 (0.10) 0.22* (0.10) 0.12 (0.09) PTD3 113 0.52 -8.59*** (2.09) 0.26 (0.14) 0.37 (0.20) 1.43** (0.50) -0.05 (0.09) 0.02 (0.08) 0.12 (0.10) 0.20* (0.10) 0.13 (0.08) PTD4 111 0.58 -12.01*** (2.02) 0.39** (0.14) 0.32* (0.16) 1.53*** (0.37) 1.28* (0.61) -0.05 (0.10) 0.01 (0.10) 0.10 (0.10) 0.17 (0.11) 0.13 (0.09) Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$ PTD1 122 0.40 -1.47*** (0.25) 0.44*** (0.12) -0.03 (0.07) 0.15 (0.10) 0.36** (0.12) 0.40* (0.20) 0.44* (0.20) 0.54** (0.18) 0.56** (0.19) PTD2 122 0.53 -4.85*** (1.27) 0.45*** (0.12) 0.48* (0.19) -0.03 (0.07) 0.13 (0.09) 0.37** (0.13) 0.36* (0.17) 0.40* (0.17) 0.50** (0.15) 0.52*** (0.15) PTD3 122 0.59 -8.30*** (2.04) 0.36** (0.13) 0.48** (0.18) 0.98* (0.39) -0.04 (0.07) 0.13 (0.10) 0.35* (0.15) 0.33 (0.17) 0.38* (0.17) 0.49** (0.16) 0.51** (0.16) PTD4 116 0.58 -9.54*** (2.74) 0.39* (0.16) 0.48** (0.18) 1.05* (0.41) 0.38 (0.68) -0.05 (0.07) 0.12 (0.11) 0.34* (0.15) 0.31 (0.18) 0.36* (0.18) 0.49** (0.19) 0.50** (0.19) Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$ PTD1 80 0.47 -1.30*** (0.20) 0.43*** (0.08) 0.11* (0.05) 0.18* (0.07) 0.32*** (0.08) 0.38* (0.15) 0.41** (0.15) PTD2 80 0.58 -3.89*** (1.11) 0.43*** (0.08) 0.37* (0.17) 0.10* (0.05) 0.17* (0.07) 0.34*** (0.09) 0.36** (0.14) 0.39** (0.13) PTD3 80 0.69 -7.95*** (1.72) 0.34*** (0.10) 0.38** (0.14) 1.13*** (0.31) 0.05 (0.06) 0.11 (0.08) 0.30** (0.10) 0.31* (0.13) 0.35** (0.12) PTD4 80 0.71 -10.38*** (2.07) 0.42*** (0.12) 0.38** (0.13) 1.26*** (0.26) 0.74 (0.50) 0.04 (0.06) 0.10 (0.08) 0.28** (0.10) 0.26 (0.13) 0.30* (0.13)
Table 5: Pooled regression results with year-specific dummy variables. We measure $t$ relative to 2006 (the start of our data collection period). Thus, $\textit{t}=1$” denotes a dummy variable for 2007, $\textit{t}=2$” denotes a dummy variable for 2008, and so on. The table contains results for multiple model specifications, named PTD1 to PTD4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
5.2 Fixed Effects Regression Results
Table 6 presents our fixed effects regression results. As before, the table contains results for multiple model specifications, now named FE1 to FE4, and multiple dependent variables as explained in Section 4. For illustration purposes, Figure 4 depicts our most basic fixed effects regression model (F1) without any control variables. A Q-Q plot of residuals is displayed in Appendix A.
Model ID $n$ $R^{2}$ (within) F (poolability) $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$
Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$
FE1 141 0.20 17.27*** 0.71*** (0.18)
FE2 141 0.20 13.74*** 0.73*** (0.17) 3.08 (10.23)
FE3 141 0.21 11.06*** 0.75*** (0.16) 3.84 (9.99) 0.82 (1.07)
FE4 132 0.25 10.44*** 0.80*** (0.19) 4.67 (11.00) 0.83 (1.20) -1.37 (1.24)
Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$
FE1 147 0.18 19.18*** 0.68** (0.21)
FE2 147 0.18 14.56*** 0.67*** (0.20) -1.28 (11.09)
FE3 147 0.18 11.03*** 0.68*** (0.19) -0.84 (10.70) 0.45 (1.12)
FE4 137 0.22 11.11*** 0.67*** (0.19) -0.15 (11.64) 0.56 (1.13) -1.74 (1.16)
Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$
FE1 135 0.16 19.67*** 0.64** (0.21)
FE2 135 0.16 15.39*** 0.64** (0.20) 1.55 (11.65)
FE3 135 0.16 11.16*** 0.68** (0.21) 2.79 (10.74) 1.27 (1.50)
FE4 128 0.17 9.37*** 0.67** (0.23) 3.35 (11.55) 0.96 (1.65) -0.99 (0.84)
Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$
FE1 125 0.02 17.74*** 0.20 (0.24)
FE2 125 0.03 14.45*** 0.20 (0.24) 5.71 (12.71)
FE3 125 0.04 10.15*** 0.22 (0.26) 6.93 (12.17) 1.54 (1.80)
FE4 121 0.06 8.91*** 0.21 (0.28) 6.80 (12.07) 1.25 (1.86) -1.31 (0.74)
Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$
FE1 113 0.02 24.66*** 0.19 (0.26)
FE2 113 0.06 21.14*** 0.18 (0.26) 8.39 (7.06)
FE3 113 0.07 14.49*** 0.19 (0.26) 9.17 (7.31) 1.39 (1.41)
FE4 111 0.09 12.20*** 0.24 (0.29) 8.40 (6.88) 1.51 (1.60) -1.01 (0.81)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
FE1 122 0.22 22.22*** 0.76** (0.26)
FE2 122 0.22 16.58*** 0.77** (0.26) 3.93 (11.31)
FE3 122 0.22 14.04*** 0.78** (0.25) 4.19 (11.18) 0.32 (1.05)
FE4 116 0.26 14.49*** 0.76** (0.25) 5.40 (11.78) 0.56 (1.10) -1.78 (1.41)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
FE1 80 0.14 17.23*** 0.53* (0.25)
FE2 80 0.18 14.09*** 0.53* (0.26) 8.68 (13.63)
FE3 80 0.23 10.68*** 0.60* (0.23) 9.38 (12.89) 3.05** (0.98)
FE4 80 0.30 10.15*** 0.58** (0.21) 10.42 (12.95) 3.05*** (0.77) -2.14 (1.72)
Table 6: Fixed effects regression results. The table contains results for multiple model specifications, named FE1 to FE4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.
All models in Table 6 have significant F-statistics for poolability, indicating that fixed effects regression models are appropriate (and that pooled regression models are not). Our $\log_{10}(STR_{(c,t)})$ coefficients are statistically significant except when we use $\log_{10}(CON_{(c,t+4)})$ or $\log_{10}(CON_{(c,t+5)})$ as the dependent variable, indicating that the relationship between STRs and convictions might disappear after three years. When significant, the fixed effects $\log_{10}(STR_{(c,t)})$ coefficients are generally substantially larger than the pooled $\log_{10}(STR_{(c,t)})$ coefficients (compare Tables 3 and 6), indicating that the relationship between STRs and convictions is affected by cross-country differences. Our largest $\log_{10}(STR_{(c,t)})$ coefficient equals $0.80$ with a standard error of $0.19$ and a $95\%$ confidence interval of $[0.42:1.18]$. While the confidence interval allows a coefficient larger than 1, our results consistently suggest a sub-linear relationship between STRs and convictions.
Fig. 4: Illustration of our most basic fixed effects regression model (FE1). $STR_{(c,t)}$ denotes the number of STRs filed in country $c$ in year $t$ per 100,000 capita. $CON_{(c,t+1:3)}$ denotes the average number of money laundering convictions per year in country $c$ up to three years after $t$ per 100,000 capita. We measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
Country $\alpha_{(c)}$ $A_{(c)}$
Spain 0.03 1.06
Italy -0.88 0.13
Germany -0.91 0.12
Belgium -1.27 0.05
Lithuania -1.34 0.05
Sweden -1.51 0.03
Portugal -1.60 0.03
United Kingdom -1.68 0.02
Czech Republic -1.69 0.02
Slovenia -1.74 0.02
Estonia -1.79 0.02
Austria -1.83 0.01
Poland -1.85 0.01
Slovak Republic -1.93 0.01
Bulgaria -1.93 0.01
Denmark -2.02 0.01
Netherlands -2.04 0.01
Finland -2.06 0.01
Croatia -2.37 0.00
Hungary -2.41 0.00
Ireland -3.31 0.00
Table 7: Estimates of country fixed effects $\alpha_{(c)}$ and $A_{(c)}$ using model FE1 and dependent variable $\log_{10}(CON_{(c,t+1:3)})$. The table should not be seen as a ranking; there are many reasons why countries might have different $\alpha_{(c)}$ and $A_{(c)}$ values. Values are impacted by our scaling.
To complement Figure 4, Table 7 presents estimates of the country-specific fixed effects $A_{(c)}$ and $\alpha_{(c)}$ estimated in model FE1 with $\log_{10}(CON_{(c,t+1:3)})$ as the dependent variable (i.e., the model depicted in Figure 4). Recall that the effects measure a country’s “scale of production;” see equations (4) and (5). Note that $\alpha_{(c)}$ is measured on a $\log_{10}$-scale (why it can be negative). The effects may reflect different legal frameworks, crime types, prosecutorial priorities, or how financial intelligence units work. In particular, the variety in the estimated fixed effects could indicate that AML regulation and enforcement is not homogenized across EU countries. We stress that the table should not be seen as a ranking. Spain clearly stands out in the table; likely an indication that it handles (or counts) money laundering cases differently. [^20] This observation makes the use of a fixed effects model more appropriate.
For robustness, we sequentially exclude any one country from our dataset and re-run all models. Regardless of which country we exclude, the $\log_{10}(STR_{(c,t)})$ coefficients remain significant when we use $\log_{10}(CON_{(c,t+1)}),\log_{10}(CON_{(c,t+2)}),\log_{10}(CON_{(c,t+3)})$, or $\log_{10}(CON_{(c,t+1:3)})$ as the dependent variable. Using $(CON_{(c,t+1:5)})$ as the dependent variable, the $\log_{10}(STR_{(c,t)})$ coefficients become insignificant in one or more models when we exclude either Austria, Finland, Germany, Lithuania, Portugal, Slovenia, or Sweden. We also consider all models with, respectively, (i) $t$ as a regressor (assuming a log-linear time trend) and (ii) time fixed effects (alongside our country-effects). For the latter, we cluster standard errors both at the unit and time level as implemented by the PanelOLS class of linearmodels [65]. Table 8 displays our results using $t$ as a regressor. All coefficients associated with $t$ are positive and significant, indicating a log-linear time trend. Meanwhile, the $\log_{10}(STR_{(c,t)})$ coefficients become much smaller and insignificant. [^21] One might suspect that the vanishing significance of the $\log_{10}(STR_{(c,t)})$ coefficients is due to “too little variation” in STRs within countries. However, Figure 4, employing a log scale, shows that there is substantial variation in STRs within countries. Also, as noted in Section 3, STRs per 100,000 capita within any given country on average (taken across all countries) increased by almost 240% between 2006 and 2014 (the median increase was approximately 140%). Most (though not all) standard errors associated with the $\log_{10}(STR_{(c,t)})$ coefficients decrease slightly or stay approximately the same when we add $t$ as a regressor (compare Tables 6 and 8). Thus, the $\log_{10}(STR_{(c,t)})$ coefficients becoming insignificant appears to be due to reduced effect sizes, not increased standard errors. This suggests a simple possibility; while both STRs and convictions have increased over time, their relationship may not be causal. Instead, an unobserved third factor, associated with time, appears to be influencing both STRs and convictions. In Section 6, we speculate that this may be related to public, political, and regulatory pressure. Note, furthermore, that all F-statistics for poolability are significant in Table 8. Thus, including $t$ as a regressor does not annul the significance of our country-specific fixed effects. Table 9 displays our results when we include time-specific fixed effects in addition to our country-specific fixed effects. Again, the $\log_{10}(STR_{(c,t)})$ coefficients generally become much smaller and insignificant (or even significantly negative). This lends credibility to the idea that an unobserved third factor, associated with time, affects both STRs and convictions.
Model ID $n$ $R^{2}$ (within) F (poolability) $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$ $t$
Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$
FET1 141 0.42 21.75*** 0.16 (0.16) 0.08*** (0.02)
FET2 141 0.42 17.60*** 0.14 (0.17) -1.79 (4.60) 0.08*** (0.02)
FET3 141 0.47 15.77*** -0.13 (0.19) -7.10 (5.68) -4.36** (1.33) 0.10*** (0.02)
FET4 132 0.50 15.21*** -0.05 (0.18) -5.86 (6.14) -4.34** (1.30) -0.98 (0.81) 0.10*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$
FET1 147 0.41 25.51*** 0.11 (0.18) 0.07*** (0.02)
FET2 147 0.43 20.39*** 0.05 (0.20) -6.32 (6.07) 0.08*** (0.02)
FET3 147 0.47 16.76*** -0.14 (0.23) -10.61 (6.98) -3.55* (1.51) 0.09*** (0.02)
FET4 137 0.48 16.26*** -0.07 (0.23) -9.44 (7.84) -3.36* (1.49) -1.06 (0.64) 0.09*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$
FET1 135 0.38 26.47*** 0.11 (0.15) 0.08*** (0.02)
FET2 135 0.39 21.13*** 0.08 (0.15) -3.42 (6.90) 0.08*** (0.02)
FET3 135 0.39 15.30*** 0.04 (0.19) -4.47 (6.96) -0.93 (2.03) 0.08*** (0.02)
FET4 128 0.39 13.04*** 0.07 (0.20) -3.47 (7.57) -1.12 (2.13) -0.85 (0.63) 0.08*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$
FET1 125 0.35 27.71*** -0.26 (0.19) 0.09*** (0.02)
FET2 125 0.35 22.66*** -0.27 (0.19) -0.74 (6.94) 0.09*** (0.02)
FET3 125 0.35 16.05*** -0.26 (0.17) -0.48 (6.67) 0.29 (2.42) 0.09*** (0.02)
FET4 121 0.38 14.89*** -0.24 (0.17) -0.75 (5.84) 0.03 (2.48) -1.60** (0.50) 0.09*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$
FET1 113 0.27 33.54*** -0.09 (0.20) 0.07*** (0.02)
FET2 113 0.27 27.73*** -0.08 (0.21) 1.91 (4.01) 0.07*** (0.02)
FET3 113 0.27 18.88*** -0.08 (0.21) 1.95 (4.51) 0.05 (1.84) 0.07*** (0.02)
FET4 111 0.33 17.17*** -0.03 (0.21) 0.15 (3.66) 0.04 (2.01) -1.62** (0.59) 0.07*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
FET1 122 0.52 33.03*** 0.18 (0.16) 0.09*** (0.02)
FET2 122 0.52 24.65*** 0.17 (0.17) -1.23 (5.49) 0.09*** (0.02)
FET3 122 0.57 23.52*** -0.04 (0.18) -5.52 (5.66) -4.26** (1.27) 0.10*** (0.02)
FET4 116 0.59 24.37*** -0.01 (0.18) -3.91 (6.22) -4.09** (1.38) -1.37 (0.92) 0.10*** (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
FET1 80 0.63 37.39*** 0.16 (0.11) 0.09*** (0.02)
FET2 80 0.63 29.42*** 0.17 (0.10) 3.46 (6.06) 0.09*** (0.02)
FET3 80 0.64 21.50*** 0.10 (0.11) 2.65 (6.25) -1.63 (1.19) 0.09*** (0.02)
FET4 80 0.71 25.04*** 0.08 (0.08) 3.68 (5.87) -1.66 (1.24) -2.24 (1.16) 0.09*** (0.01)
Table 8: Fixed effects regression results with $t$ (i.e., “time”) included as a regressor. The table contains results for multiple model specifications, named FET1 to FET4, and multiple dependent variables. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.
Model ID $n$ $R^{2}$ (within) F (poolability) $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$
Dependent variable: $\log_{10}\left(CON_{(c,t+1)}\right)$
FETF1 141 0.08 18.84*** 0.16 (0.17)
FETF2 141 0.07 15.24*** 0.14 (0.18) -1.46 (5.49)
FETF3 141 -0.09 13.88*** -0.13 (0.17) -6.90 (6.32) -4.41*** (0.85)
FETF4 132 -0.04 12.85*** -0.05 (0.16) -5.59 (6.58) -4.36*** (0.73) -1.01 (0.97)
Dependent variable: $\log_{10}\left(CON_{(c,t+2)}\right)$
FETF1 147 0.06 21.08*** 0.13 (0.18)
FETF2 147 0.04 17.04*** 0.08 (0.19) -6.11 (6.11)
FETF3 147 -0.07 14.53*** -0.11 (0.21) -10.75 (6.87) -3.75* (1.45)
FETF4 137 0.01 13.83*** -0.04 (0.23) -9.30 (7.72) -3.47* (1.55) -1.26* (0.61)
Dependent variable: $\log_{10}\left(CON_{(c,t+3)}\right)$
FETF1 135 0.06 21.62*** 0.13 (0.14)
FETF2 135 0.04 17.41*** 0.11 (0.14) -3.27 (6.96)
FETF3 135 0.01 13.22*** 0.04 (0.17) -5.22 (7.55) -1.78 (2.15)
FETF4 128 0.03 11.27*** 0.07 (0.17) -4.02 (7.99) -2.12 (2.29) -1.17 (0.81)
Dependent variable: $\log_{10}\left(CON_{(c,t+4)}\right)$
FETF1 125 -0.07 22.90*** -0.26 (0.15)
FETF2 125 -0.08 18.74*** -0.26 (0.16) -0.44 (5.93)
FETF3 125 -0.09 13.67*** -0.28* (0.11) -1.12 (6.23) -0.76 (2.35)
FETF4 121 -0.08 13.12*** -0.29* (0.13) -1.75 (4.66) -1.49 (2.10) -1.89* (0.74)
Dependent variable: $\log_{10}\left(CON_{(c,t+5)}\right)$
FETF1 113 -0.02 27.14*** -0.08 (0.20)
FETF2 113 -0.01 22.49*** -0.08 (0.21) 1.71 (5.29)
FETF3 113 -0.00 15.56*** -0.07 (0.21) 1.95 (6.08) 0.28 (1.89)
FETF4 111 -0.00 14.15*** -0.02 (0.26) -0.36 (3.95) -0.04 (2.12) -1.72* (0.82)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
FETF1 122 0.10 28.53*** 0.20 (0.19)
FETF2 122 0.09 21.67*** 0.19 (0.19) -1.12 (5.62)
FETF3 122 -0.03 21.18*** -0.03 (0.17) -6.06 (5.59) -4.59** (1.50)
FETF4 116 0.03 21.39*** 0.01 (0.15) -4.47 (6.26) -4.44** (1.62) -1.40 (0.87)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
FETF1 80 0.08 32.81*** 0.18 (0.10)
FETF2 80 0.11 26.24*** 0.19 (0.10) 3.08 (5.95)
FETF3 80 -0.03 20.45*** 0.09 (0.07) 1.63 (5.87) -2.51 (1.90)
FETF4 80 0.04 22.67*** 0.08 (0.04) 2.80 (5.52) -2.35 (1.77) -2.22 (1.55)
Table 9: Fixed effects regression results with both country- and time-specific effects. Note that standard errors are clustered both at the country and time level. The table contains results for multiple model specifications, named FETF1 to FETF4, and multiple dependent variables. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level. Note that negative $R^{2}$ (within) values can occur due to the way the statistic is computed; see the linearmodels documentation for details [65].
We stress that the relationship between STRs and convictions only disappears when we control for both country-specific and time effects. Indeed the relationship is significant in Tables 4 and 5, presenting results from models that only control for time. Furthermore, the relationship is significant in Table 6, presenting results from models that only control for country-specific effects; in fact, only controlling for country-specific effects appears to increase the size of the relationship between STRs and convictions (compare Tables 3 and 6). However, our results clearly show that both country-specific and time effects are highly relevant and should be included in an analysis of the relationship between STRs and convictions.
6 Conclusion and Discussion
Our paper investigates a simple but important question: do more STRs lead to more convictions for money laundering? To answer the question, we use data from authoritative sources such as Europol, the World Bank, the International Monetary Fund, and the European Sourcebook of Crime and Criminal Justice Statistics. Robust econometric analysis shows that the relationship between the number of STRs and convictions for money laundering in a country:
  1. 1.
    at best appears to follow a sub-linear power law and
  2. 2.
    at worst appears to be non-existent or very weak (and insignificant) when we control for cross-country differences and time effects.
Our first finding indicates that while more STRs may lead to more convictions, their marginal effect decreases with their amount. This would broadly fit with the “crying wolf” phenomenon proposed by [66]. One plausible explanation could be that prosecutorial resources (different from the number of police officers in a country) are constrained and may become “saturated.” Indeed, prosecutors and courts have finite resources. As the number of STRs increases, their capacity to process and pursue AML cases may not scale proportionally. This might cause diminishing marginal returns, where additional STRs generate fewer incremental convictions, suggesting the presence of a “saturation point.” Identifying this point would be critical to designing a more efficient AML system. Once established, the point could be used in policy design to lower the number of STRs filed while improving their quality (making it easier for law enforcement and prosecutors to use them). Alternatively, the point could also be used to justify increasing prosecutorial resources and capacity.
Our second finding implies that the relationship between STRs and convictions is driven, or at least regulated, by country-specific and temporal effects. Specifically, a log-linear time trend appears to influence both the volume of STRs filed and the number of convictions achieved in a country. We speculate that this is related to public, political, and regulatory pressure. As noted in Section 3, the average number of STRs in our dataset increased by almost 240% between 2006 and 2014. This is unlikely to be due to changes in money laundering activities. Rather, we believe it is due to increased public, political, regulatory pressure on financial institutions. At the same time, increased pressure may have caused authorities and prosecutors to prioritize money laundering cases (with the resources they have), leading to an increase in convictions not because more STRs have been filed, but because money laundering cases are given a higher priority. The situation could be a classic example of hidden or omitted variable bias with public and political pressure driving up both STRs and convictions. We also note that a number of broader developments such as the digitalization of finance and the declining use of cash in some countries might have contributed to the increase in STRs. As more financial transactions move into the formal and electronic financial system, they might prompt more STRs even when underlying behavior remains unchanged. Additionally, an increasing fragmentation of the online banking landscape might necessitate more STRs in order to unveil some money laundering operations.
We stress that our analysis focuses on the relationship between the number of STRs and convictions for money laundering in a country. We do not consider the size of convictions (i.e., if they pertain to big or small money laundering schemes or if significant sums are confiscated). We have no doubt that some STRs lead to important convictions and without insight into the context and magnitude of convictions, simple conviction counts might fail to capture important aspects of AML efforts. Our analysis also does not capture how STRs might help improve financial intelligence units’ and reporting entities’ understanding of money laundering. Furthermore, as one employee of the Danish Financial Intelligence Unit pointed out to us, STRs may be used to secure convictions for other crimes than money laundering (for example tax evasion). Our analysis does not addressed or capture this. However, in Appendix C, we consider three types of predicate crimes to money laundering: corruption, drug trafficking, and fraud. While the appendix does not rule out the possibility of a positive relationship between the number of STRs and convictions for predicate crimes, it also does not find clear, robust, and consistent evidence of such a relationship.
After the end of our STR data collection period, the EU introduced its 4th AML directive in an attempt to homogenize AML legislation and promote a risk-based approach. Our data does not allow us to evaluate whether or not this was successful. However, as discussed in Section 2, we note that the volume of STRs in the EU does not appear to have decreased after the 4th directive was introduced. We believe that this only emphasizes the importance of “quality vs. quantity considerations” in relation to STRs.
Our results suggest that financial regulators and supervisors should focus less on the number of STRs filed in a given country. [^22] Some readers might argue that STR volumes have preventive effects in and of themselves. However, we assert that money launderers do not fear STRs per se; rather, they fear the consequences STRs might bring, including asset freezes, prosecutions, and convictions. In particular, we note that regulated entities generally are prohibited from disclosing that an STR has been filed on someone (criminal or not). Thus, unless an investigation is opened, an STR might in no way affect the reported subject. Therefore, if an increase in STRs does not lead to an increase in convictions, it weakens the claim that an increase in STRs has a preventive effect. Imagine, also, a country where many STRs are filed but no one ever is convicted of money laundering: in such a country, a large volume of STRs might even help money launderers to “hide among the masses” (in line with the “crying wolf” phenomenon). Instead, they should focus more on how STRs are utilized and whether or not the “right” STRs are being filed. While our results might warrant skepticism about the effectiveness of current AML efforts and policies, we stress that we do not question the importance of such efforts and policies. We believe empirical AML research and outcome-driven compliance evaluations are highly important. We do not believe traditional AML and compliance evaluations, relying on narrative judgments, procedural checklists, and activity (rather than outcome) metrics lead to effective AML systems. Expanding empirical AML research should be a top priority for national governments and international institutions.
Appendix A Q-Q Plots of Model Residuals
Figure 6 displays a residual Q-Q plot of our pooled regression model P1 using $CON_{(c,t+1:3)}$ as the dependent variable; see Figure 1. We note deviation from the normal distribution in both the upper and lower tail. Figure 6 displays a residual Q-Q plot of our fixed effects regression model FE1 using $CON_{(c,t+1:3)}$ as the dependent variable; see Figure 4. Here, we note deviation from the normal distribution in the lower tail.
Fig. 5: Q-Q plot of residuals for pooled regression model P1 using $CON_{(c,t+1:3)}$ as the dependent variable.
Fig. 6: Q-Q plot of residuals for fixed effects model FE1 using $CON_{(c,t+1:3)}$ as the dependent variable.
Appendix B Simple Linear Models without Log-transformation
Inspired by the Cobb-Douglas production function, our main analysis employs a power law to model the relationship between STRs and convictions for money laundering. For completeness, we also include results using simple linear model formulations (i.e., without any log-transformation, simply regressing $CON_{(c,t+j)}$ on $STR_{(c,t)}$). Table 10 presents results from simple linear pooled regression models, number PL1 to PL4. Here, the $STR_{(c,t)}$ coefficients are significant (something possible even if the true relationship is a power-law). Table 11 presents results from simple linear fixed effects regression models, named FEL1 to FEL4. All models have significant F-statistics for poolability, indicating that fixed effects regression models are appropriate. Meanwhile, all $STR_{(c,t)}$ coefficients are insignificant. Thus, using simple linear models, there does not appear to be a significant relationship between STRs and convictions for money laundering when controlling for country-specific confounders.
We stress that the simple linear models are plagued by a number of problems (making us prefer the power-law formulation). Figures 7 and 8 illustrate models PL1 and FEL1 using $CON_{(c,t+1:3)}$ as the dependent variable. Here, it is easy to see how the regressions might be effected by “outlier countries.” Notably, we see very high condition numbers when fitting the models. Furthermore, as illustrated in Figures 10 and 10, the simple linear models (especially the pooled models) suffer from (highly) non-normal residuals, showing systematic curvature in residual Q–Q plots. [^23] This is not just a problem when we consider $CON_{(c,t+1:3)}$; it a general problem problem for $CON_{(c,t+j)}$. This makes confidence intervals and inference unreliable. Comparing, respectively, (i) Figures 7 and 8 with Figures 1 and 4 and (ii) Figures 10 and 10 with Figures 6 and 6, we believe a power-law model with a log-transformation (as used in our main analysis) is most appropriate.
Model ID $n$ $R^{2}$ $\alpha$ STR POP SGDP POL
Dependent variable: $CON_{(c,t+1)}$
PL1 148 0.18 6.91e-01** (2.43e-01) 1.41e-03** (5.22e-04)
PL2 148 0.32 2.82e-01 (1.65e-01) 1.50e-03** (5.61e-04) 2.06e-08 (1.14e-08)
PL3 148 0.37 -4.10e-01 (3.51e-01) 1.36e-03* (5.66e-04) 1.97e-08 (1.14e-08) 1.35e-04 (7.25e-05)
PL4 139 0.41 -2.11e+00** (6.93e-01) 1.63e-03** (5.72e-04) 1.84e-08 (1.07e-08) 1.97e-04** (6.38e-05) 4.04e-03* (1.63e-03)
Dependent variable: $CON_{(c,t+2)}$
PL1 153 0.15 8.25e-01** (2.92e-01) 1.41e-03** (5.48e-04)
PL2 153 0.29 3.69e-01 (2.03e-01) 1.53e-03** (5.85e-04) 2.22e-08 (1.26e-08)
PL3 153 0.36 -5.16e-01 (3.69e-01) 1.36e-03* (5.85e-04) 2.08e-08 (1.25e-08) 1.72e-04* (8.09e-05)
PL4 143 0.41 -2.42e+00*** (6.52e-01) 1.65e-03** (5.79e-04) 1.92e-08 (1.16e-08) 2.40e-04*** (6.62e-05) 4.55e-03** (1.70e-03)
Dependent variable: $CON_{(c,t+3)}$
PL1 141 0.13 9.04e-01** (3.11e-01) 1.40e-03** (5.36e-04)
PL2 141 0.28 4.25e-01 (2.22e-01) 1.53e-03** (5.65e-04) 2.30e-08 (1.31e-08)
PL3 141 0.36 -5.44e-01 (3.73e-01) 1.35e-03* (5.65e-04) 2.13e-08 (1.29e-08) 1.88e-04* (8.33e-05)
PL4 134 0.42 -2.43e+00*** (5.60e-01) 1.68e-03** (5.38e-04) 1.90e-08 (1.13e-08) 2.47e-04*** (6.08e-05) 4.66e-03** (1.65e-03)
Dependent variable: $CON_{(c,t+4)}$
PL1 129 0.10 1.03e+00** (3.33e-01) 1.24e-03** (4.70e-04)
PL2 129 0.24 5.30e-01* (2.57e-01) 1.38e-03** (4.73e-04) 2.36e-08 (1.36e-08)
PL3 129 0.35 -5.77e-01 (3.89e-01) 1.20e-03* (4.69e-04) 2.14e-08 (1.32e-08) 2.14e-04* (8.80e-05)
PL4 125 0.41 -2.25e+00*** (6.09e-01) 1.52e-03*** (4.21e-04) 1.83e-08 (1.11e-08) 2.58e-04*** (6.25e-05) 4.33e-03* (1.83e-03)
Dependent variable: $CON_{(c,t+5)}$
PL1 116 0.08 1.12e+00** (3.61e-01) 1.18e-03* (4.63e-04)
PL2 116 0.24 5.90e-01* (2.82e-01) 1.32e-03** (4.46e-04) 2.48e-08 (1.42e-08)
PL3 116 0.37 -6.44e-01 (4.00e-01) 1.14e-03* (4.44e-04) 2.20e-08 (1.36e-08) 2.37e-04* (9.61e-05)
PL4 114 0.45 -2.42e+00*** (6.69e-01) 1.53e-03*** (3.92e-04) 1.80e-08 (1.08e-08) 2.78e-04*** (6.81e-05) 4.67e-03* (2.09e-03)
Dependent variable: $CON_{(c,t+1:3)}$
PL1 122 0.18 7.32e-01** (2.52e-01) 1.38e-03** (4.76e-04)
PL2 122 0.32 3.27e-01 (1.77e-01) 1.49e-03** (5.10e-04) 1.99e-08 (1.15e-08)
PL3 122 0.38 -4.53e-01 (3.47e-01) 1.34e-03** (5.15e-04) 1.90e-08 (1.15e-08) 1.52e-04* (7.22e-05)
PL4 116 0.43 -2.22e+00** (6.90e-01) 1.62e-03** (5.09e-04) 1.77e-08 (1.07e-08) 2.17e-04*** (6.04e-05) 4.21e-03* (1.74e-03)
Dependent variable: $CON_{(c,t+1:5)}$
PL1 80 0.23 6.85e-01** (2.11e-01) 1.32e-03*** (3.99e-04)
PL2 80 0.35 3.45e-01* (1.64e-01) 1.42e-03** (4.35e-04) 1.66e-08 (9.67e-09)
PL3 80 0.43 -5.02e-01 (3.26e-01) 1.28e-03** (4.40e-04) 1.58e-08 (9.85e-09) 1.64e-04* (6.80e-05)
PL4 80 0.53 -2.24e+00*** (6.64e-01) 1.57e-03*** (4.34e-04) 1.52e-08 (8.73e-09) 2.28e-04*** (5.24e-05) 4.18e-03* (1.73e-03)
Table 10: Pooled regression results using a simple linear model. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
Fig. 7: Illustration of linear pooled regression model (PL1). $STR_{(c,t)}$ denotes the number of STRs filed in country $c$ in year $t$ per 100,000 capita. $CON_{(c,t+1:3)}$ denotes the average number of money laundering convictions per year in country $c$ up to three years after $t$ per 100,000 capita. We measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
Model ID $n$ $R^{2}$ (within) F (poolability) STR POP SGDP POL
Dependent variable: $CON_{(c,t+1)}$
FEL1 148 0.09 36.09*** 1.08e-03 (9.11e-04)
FEL2 148 0.10 28.70*** 1.09e-03 (9.23e-04) -9.54e-08 (1.54e-07)
FEL3 148 0.11 26.26*** 1.06e-03 (8.72e-04) -1.19e-07 (1.75e-07) -2.21e-04 (3.80e-04)
FEL4 139 0.12 21.13*** 1.07e-03 (8.81e-04) -2.34e-08 (2.02e-07) -3.53e-04 (4.07e-04) -3.15e-04 (3.55e-03)
Dependent variable: $CON_{(c,t+2)}$
FEL1 153 0.01 41.07*** 5.07e-04 (5.66e-04)
FEL2 153 0.01 33.02*** 4.98e-04 (5.66e-04) 3.92e-08 (1.43e-07)
FEL3 153 0.03 29.53*** 4.81e-04 (5.12e-04) -5.11e-09 (1.47e-07) -3.09e-04 (4.20e-04)
FEL4 143 0.11 27.17*** 4.95e-04 (5.44e-04) 5.93e-08 (2.05e-07) -3.86e-04 (3.62e-04) -4.46e-03 (3.78e-03)
Dependent variable: $CON_{(c,t+3)}$
FEL1 141 0.00 37.65*** 1.14e-04 (2.80e-04)
FEL2 141 0.02 30.87*** 9.40e-05 (2.61e-04) 2.14e-07 (2.47e-07)
FEL3 141 0.03 26.90*** 5.73e-05 (2.28e-04) 1.67e-07 (2.13e-07) -2.74e-04 (4.89e-04)
FEL4 134 0.12 25.06*** 1.52e-04 (2.72e-04) 1.70e-07 (2.46e-07) -3.38e-04 (4.57e-04) -5.25e-03 (3.45e-03)
Dependent variable: $CON_{(c,t+4)}$
FEL1 129 0.03 37.90*** -5.81e-04 (5.71e-04)
FEL2 129 0.09 33.12*** -6.03e-04 (5.77e-04) 3.87e-07 (3.49e-07)
FEL3 129 0.09 27.67*** -6.09e-04 (5.69e-04) 3.80e-07 (3.30e-07) -4.14e-05 (3.53e-04)
FEL4 125 0.17 26.64*** -3.91e-04 (3.10e-04) 3.25e-07 (3.11e-07) -7.47e-05 (3.39e-04) -5.67e-03** (2.13e-03)
Dependent variable: $CON_{(c,t+5)}$
FEL1 116 0.01 37.18*** -3.42e-04 (5.42e-04)
FEL2 116 0.12 33.90*** -3.58e-04 (5.46e-04) 4.62e-07 (3.76e-07)
FEL3 116 0.12 27.37*** -3.28e-04 (5.08e-04) 4.84e-07 (3.71e-07) 2.38e-04 (2.19e-04)
FEL4 114 0.18 25.23*** -1.23e-04 (2.72e-04) 4.42e-07 (3.79e-07) 1.32e-04 (2.67e-04) -6.05e-03** (2.06e-03)
Dependent variable: $CON_{(c,t+1:3)}$
FEL1 122 0.01 59.33*** 3.27e-04 (3.95e-04)
FEL2 122 0.02 48.20*** 3.26e-04 (3.93e-04) 5.85e-08 (1.66e-07)
FEL3 122 0.04 43.88*** 3.29e-04 (3.54e-04) 2.91e-08 (1.74e-07) -2.80e-04 (3.64e-04)
FEL4 116 0.09 36.95*** 3.20e-04 (3.51e-04) 7.21e-08 (2.16e-07) -3.62e-04 (3.63e-04) -2.25e-03 (2.75e-03)
Dependent variable: $CON_{(c,t+1:5)}$
FEL1 80 0.02 83.59*** 1.81e-04 (2.44e-04)
FEL2 80 0.03 70.36*** 1.85e-04 (2.41e-04) 1.04e-07 (1.90e-07)
FEL3 80 0.03 59.65*** 1.85e-04 (2.60e-04) 1.04e-07 (1.91e-07) -3.16e-06 (1.84e-04)
FEL4 80 0.05 48.68*** 1.63e-04 (2.56e-04) 9.95e-08 (1.98e-07) 2.46e-06 (1.92e-04) -1.68e-03 (2.22e-03)
Table 11: Fixed effects regression results using a simple linear model. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
Fig. 8: Illustration of linear fixed effects regression model (EFL1). $STR_{(c,t)}$ denotes the number of STRs filed in country $c$ in year $t$ per 100,000 capita. $CON_{(c,t+1:3)}$ denotes the average number of money laundering convictions per year in country $c$ up to three years after $t$ per 100,000 capita. We measure $t$ relative to 2006 (the start of our data collection). Thus, $\textit{t}=0$ corresponds to 2006, $\textit{t}=1$ corresponds to 2007, and so on.
Fig. 9: Q-Q plot of residuals from linear pooled regression model PL1 using $CON_{(c,t+1:3)}$ as the dependent variable.
Fig. 10: Q-Q plot of residuals from linear fixed effects regression model FEL1 using $CON_{(c,t+1:3)}$ as the dependent variable.
Appendix C Predicate Crimes to Money Laundering
Our main analysis considers the number of STRs and convictions for money laundering in a country. However, the European Sourcebook of Crime and Criminal Justice Statistics also contains statistics on some predicate crimes to money laundering, including corruption, drug trafficking and fraud. Motivated by this, we collect and model the following variables from the Sourcebook:
  • $CRP_{(c,t)}$: the number of people convicted for corruption (Sourcebook indicator code T31CO) in country $c$ in year $t$,
  • $DRT_{(c,t)}$: the number of people convicted for drug trafficking (Sourcebook indicator code T31DT) in country $c$ in year $t$, and
  • $FRD_{(c,t)}$: the number of people convicted for fraud (Sourcebook indicator code T31FR) in country $c$ in year $t$,
We collect and measure the variables analogously to $CON_{(c,t)}$ as described in Subsection 3.1. For brevity, we primarily consider fixed effects regression models (our most robust models) in this appendix. Thus, we log-transform and model each of the variables analogously to $log_{10}(CON_{(c,t+j)})$ as described in Subsection 4.2.
Table 13 displays results for convictions for corruption. Here, no $\log_{10}(STR_{(c,t)})$ coefficients are significant. If we consider simple linear models between $STR_{(c,t)}$ and $CRP_{(c,t+j)}$ (analogously to Appendix B; omitted for brevity), no $STR_{(c,t)}$ coefficients are significant. Table 13 displays results for convictions for drug trafficking. Here, some $\log_{10}(STR_{(c,t)})$ coefficients are significant for $log_{10}(DRT_{(c,t+1)})$, $log_{10}(DRT_{(c,t+4)})$, and $log_{10}(DRT_{(c,t+5)})$. Some readers might see this as evidence that STRs affect convictions for drug-trafficking. We, however, call for caution: we are considering several outcomes (i.e., convictions for corruption, drug trafficking, and fraud) and multiple time horizons without adjusting for multiplicity or time effects. If we consider simple linear models between $STR_{(c,t)}$ and $DRT_{(c,t+j)}$ (analogously to Appendix B; omitted for brevity), we have some significant $STR_{(c,t)}$ coefficients when considering $DRT_{(c,t+1)},DRT_{(c,t+3)},DRT_{(c,t+4)},$ and $DRT_{(c,t+5)}$. Finally, Table 14 contains results for convictions for fraud. Here, no $\log_{10}(STR_{(c,t)})$ coefficients are significant. If we consider simple linear models between $STR_{(c,t)}$ and $FRD_{(c,t+j)}$ (analogously to Appendix B; omitted for brevity), a single model specification has a significant $STR_{(c,t)}$ coefficient when considering $FRD_{(c,t+1:5)}$. We, again, caution against over-interpretation.
We stress that our analysis of the relationship between the number of STRs and convictions for predicate crimes is limited (especially compared to our main analysis, considering convictions for money laundering). [^24] In particular, a robust analysis of the relationship between STRs and convictions for predicate crimes would consider time effects (as our main analysis). This is outside of the scope of the present appendix (possibly warranting a study of its own). We do not rule out that there is a positive relationship between STRs and convictions for predicate crimes. However, we also fail to find robust and consistent evidence of such relationships when considering convictions for corruption, drug trafficking, and fraud. The most plausible relationship might (note our cautious wording) be between the number of STRs and convictions for drug trafficking.
Model ID $n$ $R^{2}$ (within) F (poolability) $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$
Dependent variable: $\log_{10}\left(CRP_{(c,t+1)}\right)$
FECRP1 162 0.03 85.66*** -0.15 (0.13)
FECRP2 162 0.09 86.69*** -0.17 (0.13) -6.67 (3.96)
FECRP3 162 0.10 45.64*** -0.17 (0.13) -7.96 (4.03) -1.20 (0.85)
FECRP4 147 0.16 43.36*** -0.20 (0.13) -11.13** (3.37) -1.36 (0.91) -0.03 (0.59)
Dependent variable: $\log_{10}\left(CRP_{(c,t+2)}\right)$
FECRP1 157 0.02 83.85*** -0.12 (0.15)
FECRP2 157 0.05 81.43*** -0.13 (0.15) -4.82 (3.74)
FECRP3 157 0.05 41.30*** -0.13 (0.15) -5.73 (3.91) -0.82 (0.76)
FECRP4 141 0.08 37.78*** -0.17 (0.15) -7.59 (4.02) -1.02 (0.75) 0.16 (0.68)
Dependent variable: $\log_{10}\left(CRP_{(c,t+3)}\right)$
FECRP1 139 0.00 83.20*** -0.04 (0.14)
FECRP2 139 0.00 77.36*** -0.05 (0.14) -1.58 (4.09)
FECRP3 139 0.01 37.15*** -0.05 (0.14) -2.55 (4.23) -1.02 (0.68)
FECRP4 126 0.04 33.03*** -0.08 (0.15) -2.77 (4.45) -1.12 (0.80) -1.00 (0.57)
Dependent variable: $\log_{10}\left(CRP_{(c,t+4)}\right)$
FECRP1 121 0.00 79.31*** 0.00 (0.08)
FECRP2 121 0.00 72.60*** 0.00 (0.08) -0.83 (4.06)
FECRP3 121 0.00 33.33*** -0.00 (0.08) -1.30 (3.93) -0.61 (0.89)
FECRP4 111 0.02 29.33*** 0.01 (0.09) -1.40 (3.98) -0.72 (1.07) -0.68 (0.34)
Dependent variable: $\log_{10}\left(CRP_{(c,t+5)}\right)$
FECRP1 103 0.01 74.22*** 0.12 (0.06)
FECRP2 103 0.02 67.59*** 0.12 (0.06) -2.32 (3.92)
FECRP3 103 0.02 31.48*** 0.12 (0.07) -2.15 (4.14) 0.41 (1.05)
FECRP4 96 0.03 25.68*** 0.14 (0.08) -2.88 (4.30) 0.60 (1.05) -0.22 (0.68)
Dependent variable: $\log_{10}\left(CRP_{(c,t+1:3)}\right)$
FECRP1 136 0.04 152.87*** -0.14 (0.16)
FECRP2 136 0.09 153.66*** -0.14 (0.16) -5.13 (4.43)
FECRP3 136 0.11 77.32*** -0.15 (0.15) -6.15 (4.43) -1.01 (0.69)
FECRP4 124 0.19 74.31*** -0.19 (0.16) -7.85 (4.12) -1.13 (0.81) -0.61 (0.50)
Dependent variable: $\log_{10}\left(CRP_{(c,t+1:5)}\right)$
FECRP1 96 0.02 259.31*** -0.08 (0.13)
FECRP2 96 0.05 251.74*** -0.07 (0.13) -3.58 (4.83)
FECRP3 96 0.05 113.72*** -0.07 (0.13) -3.63 (4.90) -0.07 (0.65)
FECRP4 90 0.09 98.08*** -0.10 (0.13) -4.18 (5.05) -0.21 (0.79) -0.55 (0.61)
Table 12: Fixed effects regression results considering convictions for corruption. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.
Model ID $n$ $R^{2}$ (within) F (poolability) $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$
Dependent variable: $\log_{10}\left(DRT_{(c,t+1)}\right)$
FEDRT1 124 0.11 84.09*** 0.13* (0.06)
FEDRT2 124 0.12 76.88*** 0.13* (0.06) -1.40 (3.16)
FEDRT3 124 0.16 46.33*** 0.15* (0.06) -0.21 (2.83) 1.00* (0.44)
FEDRT4 115 0.19 29.61*** 0.15* (0.06) -0.36 (2.64) 1.05* (0.49) -0.51 (0.58)
Dependent variable: $\log_{10}\left(DRT_{(c,t+2)}\right)$
FEDRT1 123 0.05 87.71*** 0.08 (0.05)
FEDRT2 123 0.05 77.89*** 0.08 (0.05) 0.22 (3.14)
FEDRT3 123 0.08 49.02*** 0.10 (0.05) 1.13 (2.95) 0.80* (0.38)
FEDRT4 112 0.14 36.45*** 0.08 (0.05) 0.94 (2.57) 0.95* (0.43) -0.79 (0.70)
Dependent variable: $\log_{10}\left(DRT_{(c,t+3)}\right)$
FEDRT1 111 0.02 89.29*** 0.05 (0.03)
FEDRT2 111 0.03 79.53*** 0.05 (0.03) 0.99 (3.33)
FEDRT3 111 0.07 48.53*** 0.07 (0.04) 1.81 (3.08) 0.96* (0.44)
FEDRT4 101 0.07 31.82*** 0.06 (0.04) 1.30 (2.90) 0.97* (0.48) -0.26 (0.47)
Dependent variable: $\log_{10}\left(DRT_{(c,t+4)}\right)$
FEDRT1 98 0.03 108.64*** 0.06 (0.03)
FEDRT2 98 0.05 97.70*** 0.06 (0.04) 1.50 (3.18)
FEDRT3 98 0.07 55.74*** 0.07* (0.04) 1.88 (3.10) 0.66 (0.53)
FEDRT4 89 0.05 33.50*** 0.06* (0.03) 1.35 (3.09) 0.66 (0.52) 0.09 (0.30)
Dependent variable: $\log_{10}\left(DRT_{(c,t+5)}\right)$
FEDRT1 85 0.08 135.58*** 0.09* (0.04)
FEDRT2 85 0.11 124.77*** 0.09* (0.04) 2.33 (3.00)
FEDRT3 85 0.11 72.35*** 0.09* (0.04) 2.36 (2.95) 0.30 (0.63)
FEDRT4 78 0.09 43.72*** 0.08* (0.03) 1.74 (3.08) 0.30 (0.66) 0.16 (0.33)
Dependent variable: $\log_{10}\left(DRT_{(c,t+1:3)}\right)$
FEDRT1 99 0.07 103.15*** 0.09 (0.05)
FEDRT2 99 0.07 87.53*** 0.09 (0.05) -0.12 (3.58)
FEDRT3 99 0.11 49.71*** 0.10 (0.06) 0.86 (3.40) 0.89** (0.32)
FEDRT4 93 0.20 34.65*** 0.09 (0.06) 0.79 (2.91) 0.98* (0.39) -0.86 (0.72)
Dependent variable: $\log_{10}\left(DRT_{(c,t+1:5)}\right)$
FEDRT1 66 0.03 137.57*** 0.05 (0.05)
FEDRT2 66 0.03 106.18*** 0.05 (0.05) 0.64 (4.14)
FEDRT3 66 0.13 53.78*** 0.07 (0.05) 1.20 (4.03) 1.05*** (0.29)
FEDRT4 64 0.28 38.79*** 0.06 (0.06) 0.65 (3.29) 0.75* (0.31) -1.31 (0.92)
Table 13: Fixed effects regression results considering convictions for drug trafficking. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.
Model ID $n$ $R^{2}$ (within) F (poolability) $\log_{10}\left(STR_{(c,t)}\right)$ $\log_{10}\left(POP_{(c,t)}\right)$ $\log_{10}\left(SGDP_{(c,t)}\right)$ $\log_{10}\left(POL_{(c,t)}\right)$
Dependent variable: $\log_{10}\left(FRD_{(c,t+1)}\right)$
FEFRD1 181 0.00 38.49*** -0.03 (0.13)
FEFRD2 181 0.10 42.06*** -0.03 (0.12) -6.92* (3.19)
FEFRD3 181 0.11 42.11*** -0.04 (0.12) -7.84** (3.00) -0.82 (0.86)
FEFRD4 163 0.18 37.26*** 0.05 (0.08) -8.30** (2.59) -1.07 (0.61) -1.00 (0.72)
Dependent variable: $\log_{10}\left(FRD_{(c,t+2)}\right)$
FEFRD1 179 0.00 38.51*** -0.02 (0.11)
FEFRD2 179 0.08 42.12*** -0.01 (0.11) -6.40* (2.97)
FEFRD3 179 0.11 43.29*** -0.03 (0.10) -7.97** (2.94) -1.43 (0.75)
FEFRD4 159 0.18 35.83*** 0.02 (0.09) -7.85** (2.69) -1.71** (0.64) -1.14 (0.61)
Dependent variable: $\log_{10}\left(FRD_{(c,t+3)}\right)$
FEFRD1 158 0.00 37.84*** -0.01 (0.08)
FEFRD2 158 0.04 39.35*** -0.01 (0.08) -4.73 (2.98)
FEFRD3 158 0.07 40.19*** -0.02 (0.08) -6.03 (3.24) -1.28 (0.74)
FEFRD4 142 0.10 31.13*** 0.00 (0.08) -5.71 (3.09) -1.41* (0.71) -0.82 (0.44)
Dependent variable: $\log_{10}\left(FRD_{(c,t+4)}\right)$
FEFRD1 138 0.01 47.23*** 0.05 (0.05)
FEFRD2 138 0.02 47.41*** 0.05 (0.06) -2.90 (2.37)
FEFRD3 138 0.04 47.90*** 0.05 (0.06) -3.71 (2.55) -0.93 (0.71)
FEFRD4 126 0.07 38.85*** 0.07 (0.06) -3.50 (2.36) -0.95 (0.70) -0.65* (0.29)
Dependent variable: $\log_{10}\left(FRD_{(c,t+5)}\right)$
FEFRD1 120 0.01 52.25*** 0.06 (0.06)
FEFRD2 120 0.01 51.20*** 0.06 (0.06) -1.23 (1.71)
FEFRD3 120 0.02 51.05*** 0.06 (0.06) -0.87 (1.89) 0.56 (0.89)
FEFRD4 112 0.10 46.75*** 0.10 (0.05) -1.13 (1.64) 0.51 (0.89) -1.00*** (0.28)
Dependent variable: $\log_{10}\left(FRD_{(c,t+1:3)}\right)$
FEFRD1 146 0.00 49.37*** 0.01 (0.09)
FEFRD2 146 0.07 52.87*** 0.00 (0.09) -5.52 (3.52)
FEFRD3 146 0.10 54.23*** -0.00 (0.09) -6.84 (3.56) -1.24 (0.74)
FEFRD4 132 0.22 47.21*** 0.02 (0.08) -6.52* (3.00) -1.34* (0.60) -1.26 (0.64)
Dependent variable: $\log_{10}\left(FRD_{(c,t+1:5)}\right)$
FEFRD1 99 0.02 100.30*** 0.06 (0.05)
FEFRD2 99 0.11 108.65*** 0.06 (0.06) -4.93 (3.53)
FEFRD3 99 0.17 115.32*** 0.06 (0.05) -5.81 (3.51) -1.19 (0.67)
FEFRD4 93 0.34 117.29*** 0.04 (0.07) -5.30* (2.54) -1.38* (0.64) -1.41* (0.56)
Table 14: Fixed effects regression results considering convictions for fraud. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level.